A geo-stationary satellite orbits around the earth in a circular orbit of radius 36,000 km. Then, the time period of a spy satellite orbiting a few hundred km above the earth’s surface (Rearth = 6,400km) will approximately be
(a) ½ hr
(b) 1 hr
(c) 2 hr
(d) 4 hr

Grade:6

2 Answers

Deepak Patra
9 years ago

(c) Note : A satellite revolving near the earth’s surface has a time period of 84.6 min.

We know that as the height increases, the time period increases. Thus the time period of the spy satellite should be slightly greater than 84.6 minutes.

T = 2 hr
Riya
4 years ago
we know that a time period of geostationary satellite is 24 hours’
T^2 is directly proportional to R^3
T2/T1=(R2/R1)^3/2
 
T2/24=(6400/36000)^3/2
T2=24 x (6400/36000)^3/2
 
T2= 2 hours
 
hence c is the correct option.
 
Hence c is the correct answer.